# An Effective Positivstellensatz over the Rational Numbers for Finite Semialgebraic Sets

### v1

@@ 2.1 Notation and preliminaries
[…]
@@
2.2 The case when $S=V_{\mathbb{R}}(I)$ for a radical zero-dimensional ideal $I$
@@
We begin by considering in next lemma the case where $V_{\mathbb{R}}(I)\neq\emptyset$.
@@
2.3 The case $S=V_{\mathbb{R}}(I)$ for an arbitrary zero-dimensional ideal $I$
@@
Define $J:=\sqrt{I}$. Since $\mathbb{Q}[\mathbf{x}]/J\simeq(\mathbb{Q}[\mathbf{x}]/I)\big{/}(J/I)$, we can assume that a monomial basis of $\mathbb{Q}[\mathbf{x}]/J$ is included in the monomial basis $B$ of $\mathbb{Q}[\mathbf{x}]/I$. Then, by Theorem 2.1, since #$\#(V_{\mathbb{C}}(J))=D$, there exists $\omega_{k}\in\mathbb{Q}_{>0}$, $\theta_{k}\in\langle B\rangle_{\mathbb{Q}}$, $1\leq k\leq D$, where we can assume $\theta_{1}(\zeta)\neq 0$ for all $\zeta\in V_{\mathbb{C}}$, such that
@@ 3 SoS representation for $f$ nonnegative on $S$
[…]
@@
[…]
@@
Since $(I:f)+(f)=(1)$, we can find $a\in\mathbb{Q}[\mathbf{x}]$ and $b\in(I:f)$ such that $b+a\,f=1$.
@@
[…]
@@
[…]
@@ 4.1 Bounds on the values of a polynomial at the roots
Let $J=(h_{1},\dots,h_{s+1})$ be a zero-dimensional ideal defined by polynomials $h_{1},\dots,h_{s+1}\in\mathbb{Z}[\mathbf{x}]$ with $\deg(h_{j})\leq d$ for $1\leq j\leq s$, $e=\max\{d,\deg(h_{s+1})\}$ and $\mathfrak{h}(h_{j})\leq\tau$ for $11\leq j\leq s+$ 1. Then
@@
$-{\mathcal{C}}(n)\,d^{n-1}d_{p}(d+\tau)\leq\sum_{\zeta\in Z}\mathfrak{h}(p(\zeta))\leq{\mathcal{C}}(n)\,d^{n-1}d_{p}(d+\tau)$
@@ 4.2 Reduction in the quotient algebra $\mathbb{Q}[\mathbf{x}]/I$
Let $J=(h_{1},\dots,h_{s},h_{s+1})$ be a zero-dimensional radical ideal defined by polynomials $h_{1},\dots,h_{s},h_{s+1}\in\mathbb{Z}[\mathbf{x}]$ with $\deg(h_{j})\leq d$ for $1\leq j\leq s$, $e=\max\{d,\deg(h_{s+1})\}$ and $\mathfrak{h}(h_{j})\leq\tau$ for $11\leq j\leq s+$ 1. Then there exists $w_{0},w_{1},\ldots,w_{n}\in\mathbb{Z}[t]$ with $\deg(w_{0})=D:=\dim\mathbb{Q}[\mathbf{x}]/J$, $\deg(w_{i})<D$ for $1\leq i\leq n$ such that
@@
is homogeneous of degree $D$ and $-\frac{\partial W_{0}(t,U)}{\partial U_{i}}\in\mathbb{Z}[t,U]$ is homogeneous of degree $D-1$, and therefore each of the (integer) coefficients of $w_{0}(t)$ has absolute value bounded by $2^{n+D}2^{\mathfrak{h}(W_{0})}D^{\prime D}$, while each of the integer coefficients of $w_{i}(t)$, $1\leq i\leq n$, has absolute value bounded by $2^{n+D-1}2^{\mathfrak{h}(\frac{\partial W_{0}(U,t)}{\partial U_{i}})}D^{\prime D-1}$. This adds a factor $2\,n\,d^{n}\log(d)$ that we can summarize in
@@
To compute the vector $(\lambda_{1},\ldots,\lambda_{D})$ and all coefficients of all polynomials $p_{j}$, we solve a linear system $\mathbf{A}\mathbf{y}=\mathbf{b}$ where $\mathbf{A}\in\mathbb{Z}^{M\times N}$ which non-zero entries are coefficients of the $b_{i}$’s and the polynomials $h_{j}$, so that $\mathfrak{h}(\mathbf{A})\leq\tau$, and $\mathbf{b}\in\mathbb{Z}^{M}$ is the vector of coefficients of $p$, so that $\mathfrak{h}(\mathbf{b})\leq\tau_{p}$. Since the number of coefficients of $p$ is at most $d_{p}^{n}$ and we can assume that $\mathbf{A}$ is of full rank $M$, we have $M\leq d_{p}^{n}$. By applying Lemma A.1 we obtain $\nu\in\mathbb{Z}$ and the $\lambda_{i}\in\mathbb{Z}$ and the coefficients of the m polynomials $p_{k}\in\mathbb{Z}[\mathbf{x}]$ with
@@ 4.3 Bounds on the singular values of the Vandermonde matrix
by Lemma 4.1. This implies that each of the interpolant $\varphi_{\zeta}$ of degree $D-1$ satisfies
@@ 5.1 The case when $S=V_{\mathbb{R}}(I)$
$21<\lambda_{\zeta}-|\Re({p}(\zeta))|<\lambda_{\zeta}+\Re({p}(\zeta))<2\,|{p}(\zeta)|+$ 2
@@
By Remark 5.4 we now round $\widetilde{Q}\in S^{D}(\mathbb{R})$ to $2^{-N}Q$ with $Q\in S^{D}(\mathbb{Z})$ and $N\in\mathbb{N}$ with $\mathfrak{h}(N)={\mathcal{C}}(n)\,d^{n-1}\delta(d+\tau)+{\mathcal{C}}(1)\,\eta_{p}$ so that $\,\textup{dist}\,(\widetilde{Q},2^{-N}Q)<\sigma_{\min}(\widetilde{Q})$ and
@@ 5.2 The case when $f>0$ on $S$ for $S\subset V$
[…]
@@ 6 Height bounds for $I$ radical and $f$ nonnegative on $S$
[…]
@@
[…]
@@ 7 Algorithm and examples
To find such a decomposition, we solve a Semi-Definite Program that maximizes the smallest eigenvalue of $Q_{0}$ in (20). LABEL:lem:positive_definite implies that, when considering monomials of sufficiently high degree, this smallest eigenvalue will be strictly positive. Then we round, to a correct precision and taking into account the smallest eigenvalue of $Q^{*}_{0}$, the approximate optimal matrices $Q_{i}^{*}$ and polynomials $p_{j}^{*}$, to obtain an exact SoS representation for $f$.
@@
[…]
@@
  • $M_{i}\leq\widehat{d}^{n}$ (resp. $M^{\prime}_{j}\leq\widehat{d}^{n}$) (resp. $N\leq\widehat{d}^{n}$) is the number of monomials of degree $<\widehat{d}-\deg(g_{i})$ (resp. $<\widehat{d}-\deg(h_{j})$ with $g_{0}=1$) (resp. $\widehat{d}$),
@@
with G $=[g_{1}]$, H $=[h_{1},h_{2}]$, X $=[x,y]$ and a relaxation order $r=2$. We obtain the exact decomposition
@@
[…]
@@
Since $2x=ax^{2}+b\,x=a\,x^{2}-h_{1}-h_{2}$, we deduce a SoS representation of $f=x$:
@@ A.1 Solving linear systems
Let $\mathbf{A}\in\mathbb{Z}^{M\times N}$ be such that $M\leq N$ and $\mathbf{b}\in\mathbb{Z}^{M}$ with $\mathfrak{h}(\mathbf{A})\leq\tau_{A}$ and $\mathfrak{h}(\mathbf{b})\leq\tau_{b}$. If the linear system $\mathbf{A}\,\mathbf{y}=\mathbf{b}$ admits a solution, then one of these solutions can be written by Cramer’s rule as $\mathbf{c}=\widehat{\mathbf{c}}/\nu$ with $\widehat{\mathbf{c}}\in\mathbb{Z}^{D}$ and $\nu\in\mathbb{N}$ satisfying by Hadamard’s inequality
@@ A.2 Projection on a linear space
=$\mathcal{L}=\{\mathbf{y}\in\mathbb{R}^{N}\,:\,\mathbf{A}\mathbf{y}=\mathbf{b}\}.$
@@ A.3 Weighted square-root-free Cholesky decomposition
where $Q^{(k-1)}(X_{k},\ldots,X_{M})=(q_{i,j}^{(k-1)})$ with $q_{i,j}^{(k-1)}=\frac{1}{\Delta_{k-1}}\Delta^{(k-1)}_{i,j}$. By the same construction as above applied to $Q^{(k-1)}$ we obtain a decomposition at order $k$ from the decomposition (23) or order $k-1$, with $Q^{(k)}=(q^{(k)}_{i,j})$ such that
@@
This shows in particular that if $Q$ is a quadratic form such as $\Delta_{k}\neq 0$, then its signature is $(M-s,s)$ where $s$ is the number of sign changes of the sequence $[\Delta_{0}=1,\Delta_{1},\ldots,\Delta_{M}]$. This is a well-known result, probably due to Sylvester.

### v2

@@ 2.1 Notation and preliminaries
    • $u_{\overline{\zeta}}=\overline{u_{\zeta}}$, and in particular $u_{\xi}\in\mathbb{R}[\mathbf{x}]$ if $\xi\in V_{\mathbb{R}}$;
@@
2.2 The case when $S=V_{\mathbb{R}}(I)$ for a radical zero-dimensional ideal II
@@
We begin by considering in the next lemma the case where $V_{\mathbb{R}}(I)\neq\emptyset$.
@@
2.3 The case $S=V_{\mathbb{R}}(I)$ for an arbitrary zero-dimensional ideal II
@@
Define $J:=\sqrt{I}$. Since $\mathbb{Q}[\mathbf{x}]/J\simeq(\mathbb{Q}[\mathbf{x}]/I)\big{/}(J/I)$, we can assume that a monomial basis of $\mathbb{Q}[\mathbf{x}]/J$ is included in the monomial basis $B$ of $\mathbb{Q}[\mathbf{x}]/I$. Then, by Theorem 2.1, since $(V_{\mathbb{C}}(J))=D$, there exists $\omega_{k}\in\mathbb{Q}_{>0}$, $\theta_{k}\in\langle B\rangle_{\mathbb{Q}}$, $1\leq k\leq D$, where we can assume $\theta_{1}(\zeta)\neq 0$ for all $\zeta\in V_{\mathbb{C}}$, such that
@@ 3 SoS representation for $f$ nonnegative on $S$
Here we consider Problem 2 stated in Section 1, whose answer is summarized in A, when $f$ is nonnegative on $S$ and $(I:f)+(f)=(1)$.
@@
If $f\in\mathbb{Q}[\mathbf{x}]$ is such that $f\geq 0$ on $S$ and $(I:f)+(f)=(1)$, then there exist $\omega_{i,k}\in\mathbb{Q}_{\geq 0}$, $q_{i,k}\in\langle B\rangle_{\mathbb{Q}}$ and $p_{j}\in\mathbb{Q}[\mathbf{x}]$ for $0\leq i\leq r$, $1\leq k\leq D$ and $1\leq j\leq s$ such that
@@
Since $(I:f)+(f)=(1)$, we can find $a\in\mathbb{Q}[\mathbf{x}]$ and $b\in(I:f)$ such that b+af=1b+a\,f=1.
@@
In [KMS23] the authors showed that for the univariate case, the condition $h/\gcd(f,h)$ and $f$ being relatively prime is a sufficient condition in order to guarantee that if $f$ is a polynomial such that $f(\xi)\geq 0$ for all $\xi\in V_{\mathbb{R}}(I)$, then $f\in Q_{\mathbb{Q}}$ (here $I=(h)$). The condition $(I:f)+(f)=(1)$ of Theorem 3.2 is equivalent to that condition in the univariate case.
@@
A direct corollary of this theorem is the characterisation of positive polynomials on $S$, given in B. Indeed if $I$ is radical, we have $(I:f)+(f)=(1)$ since $V_{\mathbb{C}}(I:f)\cap V_{\mathbb{C}}(f)=\emptyset$, and by Theorem 3.2 any polynomial nonnegative on $S$ has the desired SoS representation.
@@ 4.1 Bounds on the values of a polynomial at the roots
Let $J=(h_{1},\dots,h_{s+1})$ be a zero-dimensional ideal defined by polynomials $h_{1},\dots,h_{s+1}\in\mathbb{Z}[\mathbf{x}]$ with $\deg(h_{j})\leq d$ for $1\leq j\leq s$, $e=\max\{d,\deg(h_{s+1})\}$ and $\mathfrak{h}(h_{j})\leq\tau$ for $1\leq j\leq s+1$. Then
@@
${\mathcal{C}}(n)\,d^{n-1}d_{p}(d+\tau)\leq\sum_{\zeta\in Z}\mathfrak{h}(p(\zeta))\leq{\mathcal{C}}(n)\,d^{n-1}d_{p}(d+\tau)$
@@ 4.2 Reduction in the quotient algebra $\mathbb{Q}[\mathbf{x}]/I$
Let $J=(h_{1},\dots,h_{s},h_{s+1})$ be a zero-dimensional radical ideal defined by polynomials $h_{1},\dots,h_{s},h_{s+1}\in\mathbb{Z}[\mathbf{x}]$ with $\deg(h_{j})\leq d$ for $1\leq j\leq s$, $e=\max\{d,\deg(h_{s+1})\}$ and $\mathfrak{h}(h_{j})\leq\tau$ for $1\leq j\leq s+1$. Then there exists $w_{0},w_{1},\ldots,w_{n}\in\mathbb{Z}[t]$ with $\deg(w_{0})=D:=\dim\mathbb{Q}[\mathbf{x}]/J$, $\deg(w_{i})<D$ for $1\leq i\leq n$ such that
@@
is homogeneous of degree $D$ and $\frac{\partial W_{0}(t,U)}{\partial U_{i}}\in\mathbb{Z}[t,U]$ is homogeneous of degree D1D-1, and therefore each of the (integer) coefficients of $w_{0}(t)$ has absolute value bounded by $2^{n+D}2^{\mathfrak{h}(W_{0})}D^{\prime D}$, while each of the integer coefficients of $w_{i}(t)$, $1\leq i\leq n$, has absolute value bounded by $2^{n+D-1}2^{\mathfrak{h}(\frac{\partial W_{0}(U,t)}{\partial U_{i}})}D^{\prime D-1}$. This adds a factor $2\,n\,d^{n}\log(d)$ that we can summarize in
@@
To compute the vector $(\lambda_{1},\ldots,\lambda_{D})$ and all coefficients of all polynomials $p_{j}$, we solve a linear system $\mathbf{A}\mathbf{y}=\mathbf{b}$ where $\mathbf{A}\in\mathbb{Z}^{M\times N}$ which non-zero entries are coefficients of the $b_{i}$’s and the polynomials $h_{j}$, so that $\mathfrak{h}(\mathbf{A})\leq\tau$, and $\mathbf{b}\in\mathbb{Z}^{M}$ is the vector of coefficients of $p$, so that $\mathfrak{h}(\mathbf{b})\leq\tau_{p}$. Since the number of coefficients of $p$ is at most $d_{p}^{n}$ and we can assume that $\mathbf{A}$ is of full rank $M$, we have $M\leq d_{p}^{n}$. By applying Lemma A.1 we obtain $\nu\in\mathbb{Z}$ and the $\lambda_{i}\in\mathbb{Z}$ and the coefficients of the polynomials $p_{j}\in\mathbb{Z}[\mathbf{x}]$ with
@@ 4.3 Bounds on the singular values of the Vandermonde matrix
by Lemma 4.1. This implies that each of the interpolant $\varphi_{\zeta}$ of degree D1D-1 satisfies
@@ 5.1 The case when $S=V_{\mathbb{R}}(I)$
$1<\lambda_{\zeta}-|\Re({p}(\zeta))|<\lambda_{\zeta}+\Re({p}(\zeta))<2\,|{p}(\zeta)|+2$
@@
[…]
@@ 5.2 The case when $f>0$ on $S$ for $S\subset V$
  • $\mathfrak{h}(\omega_{i,k})\leq{\mathcal{C}}(n)\,d^{n-1}d_{f}(d+\tau)$ for $1\leq i\leq r,\ 1\leq k\leq D$;
  • $\mathfrak{h}(q_{i,k})\leq{\mathcal{C}}(n;d)\,d^{n-1}(d^{n}\delta+d_{f})(d+\tau)$ for $0\leq i\leq r,\ 1\leq k\leq D$.
@@ 6 Height bounds for $I$ radical and $f$ nonnegative on $S$
Since $I$ is radical, we have $(I:f)+(f)=(1)$. Hereafter, we use this identity to reduce to the cases treated in Section 5.
@@
Let $B=\{\,b_{1},\dots,b_{D}\,\}$ be the basis of $\mathbb{Q}[\mathbf{x}]\big{/}I$. We first analyze polynomials $\widetilde{a}\in\langle B\rangle_{\mathbb{Q}}$ and $\widetilde{b}\in(I:f)\cap\langle B\rangle_{\mathbb{Q}}$ such that $\widetilde{a}\,f+\widetilde{b}\equiv 1\ \,\textup{mod}\,I$ (we note that such $\widetilde{a}$ and $\widetilde{b}$ exist since $(I:f)+(f)=(1)$).
@@ 7 Algorithm and examples
To find such a decomposition, we solve a Semi-Definite Program that maximizes the smallest eigenvalue of $Q_{0}$ in (20). Lemma 7.1 implies that, when considering monomials of sufficiently high degree, this smallest eigenvalue will be strictly positive. Then we round, to a correct precision and taking into account the smallest eigenvalue of $Q^{*}_{0}$, the approximate optimal matrices $Q_{i}^{*}$ and polynomials $p_{j}^{*}$, to obtain an exact SoS representation for $f$.
@@
Output: $f=\sum_{k}\widehat{\omega}_{0,k}\widehat{q}_{0,k}^{\,2}+\sum_{i}\big{(}\sum_{k}\widehat{\omega}_{i,k}\widehat{q}_{i,k}^{\,2}\big{)}g_{i}+\sum_{j}\widehat{p}_{j}\,h_{j}$ with $\widehat{\omega}_{0,k}\in\mathbb{Q}_{>0}$, $\widehat{\omega}_{i,k}\in\mathbb{Q}_{\geq 0}$,  $\widehat{q}_{i,k}\in\mathbb{Q}[\mathbf{x}]_{\ell_{i}}$, $\widehat{p}_{j}\in\mathbb{R}[\mathbf{x}]_{\max\{2\ell,\deg(g_{i})+2\ell_{i}\}-\deg(h_{k})}$.
@@
  • $M_{i}\leq\widehat{d}^{n}$ (resp. $M^{\prime}_{j}\leq\widehat{d}^{n}$) (resp. $N\leq\widehat{d}^{n}$) is the number of monomials of degree $\widehat{d}-\deg(g_{i})$ (resp. $\widehat{d}-\deg(h_{j})$ with $g_{0}=1$) (resp. $\widehat{d}$),
@@
with G $g_{1}]$, H $h_{1},h_{2}]$, X =[x,y]=[x,y] and a relaxation order $r=2$. We obtain the exact decomposition
@@
Let $f=x\in\mathbb{Q}[x,y]$, which is nonnegative on $S$. We can verify that $(I:f)+(f)=(1)$ and that ax+b=2a\,x+b=2, for $a=1+x$ and $b=2-x-x^{2}$, where $b\,x=-h_{1}-h_{2}\in I$.
@@
Since $2x=ax^{2}+b\,x=a\,x^{2}-h_{1}-h_{2}$, we deduce an SoS representation of $f=x$:
@@ Acknowledgments
Finally, the authors want to thank the referee for the helpful comments and for suggesting the reference [HQ25].
@@ A.1 Solving linear systems
Let $\mathbf{A}\in\mathbb{Z}^{M\times N}$ be such that $M\leq N$ and $\mathbf{b}\in\mathbb{Z}^{M}$ with $\mathfrak{h}(\mathbf{A})\leq\tau_{A}$ and $\mathfrak{h}(\mathbf{b})\leq\tau_{b}$. If the linear system $\mathbf{A}\,\mathbf{y}=\mathbf{b}$ admits a solution, then one of these solutions can be written by Cramer’s rule as $\mathbf{c}=\widehat{\mathbf{c}}/\nu$ with $\widehat{\mathbf{c}}\in\mathbb{Z}^{N}$ and $\nu\in\mathbb{N}$ satisfying by Hadamard’s inequality
@@ A.2 Projection on a linear space
$\mathcal{L}=\{\mathbf{y}\in\mathbb{R}^{N}\,:\,\mathbf{A}\mathbf{y}=\mathbf{b}\}.$
@@ A.3 Weighted square-root-free Cholesky decomposition
where $Q^{(k-1)}(X_{k},\ldots,X_{M})=(q_{i,j}^{(k-1)})$ with $q_{i,j}^{(k-1)}=\frac{1}{\Delta_{k-1}}\Delta^{(k-1)}_{i,j}$. By the same construction as above applied to $Q^{(k-1)}$ we obtain a decomposition at order $k$ from the decomposition (23) or order k1k-1, with $Q^{(k)}=(q^{(k)}_{i,j})$ such that
@@
This shows in particular that if $Q$ is a quadratic form such as $\Delta_{k}\neq 0$, then its signature is (Ms,s)(M-s,s) where $s$ is the number of sign changes of the sequence $\Delta_{0}=1,\Delta_{1},\ldots,\Delta_{M}]$. This is a well-known result, probably due to Sylvester.